# Driving Scene Synthesis on Free-form Trajectories with Generative Prior

### v1

@@ 4.2 Comparison with state of the art
Results on generated videos
@@
Another intriguing application of the proposed method is transforming AI-generated video into a re-enactable driving world. To quantitatively evaluate this capability, we deploy DriveX and the reconstruction-only baseline [35] on 12 driving videos generated by [7]. The FID of videos rendered across different novel trajectories is reported in Tab. 9. The results indicate that the realism of rendered videos is well preserved within a certain range of shifted length. Compared to the reconstruction-only baseline, our DriveX achieves 4.1% and 2.6% FID reduction in shifting length of ±$\pm 1m$ and ±$\pm 2m$ respectively. To obtain real world distance, we align the predicted depth and LiDAR depth of the first frame by the least square estimation. These advancements are also clearly illustrated in Fig. 4. As can be seen, the baseline method exhibits significant degradation in views that are far from the recorded trajectory.
4.3 Ablation study
[…]
PVG [4] 0.0200 0.5401 118.41
PVG [4] 0.0819 0.6188 101.04
StreetGaussian [35] 0.1217 0.6124 103.08
StreetGaussian [35] 0.2092 0.6257 90.30
Table 4: Ablation study on different Gaussian model. Unless stated otherwise, all ablation experiments are measured on views at the novel trajectory shifted $3m$ from the recorded trajectory.
[…]
mask all 0.0899 0.5781 107.58
w/o mask 0.1106 0.5965 90.49
w/ mask 0.2092 0.6257 90.30
Table 5: Effects of unreliability mask.
[…]
500 0.1070 0.5954 86.56 4.8×4.8$\times$4.8 ×
2,000 0.2092 0.6257 90.30 1.7×1.7$\times$1.7 ×
5,000 0.1104 0.5992 105.25 1.4×1.4$\times$1.4 ×
Inf 0.0945 0.5845 95.08 1.0×1.0$\times$1.0 ×
Table 6: Ablations on buffer intervals. Inf denotes the novel trajectory supervision only generated once time at iteration $T_{0}$.
(a) Ablation on unreliability mask (b) Ablation on iterative refinement
Figure 5: Ablations on (a) unreliability mask and (b) iterative refinement. Fig.(a) shows the refined results without (middle left) and with (middle right) unreliability mask (top right). The final rendered results exhibit significant artifacts without mask (bottom left). Fig.(b) shows the rendered (top) and restored results at iteration $T_{0}$ (middle left) and $T$ (middle right). It can be seen that single-time refinement can not completely eliminate artifacts, resulting in inferior final rendered results compared to iterative refinement (bottom). $T_{0}$ and $T$ indicate the first and last iteration after warm up, respectively.
Different Gaussian model
Although we adopt StreetGaussian [35] as the default Gaussian counterpart in previous experiments, our method is also capable of synergy with different choices. To demonstrate the versatility of DriveX, we evaluate on a dynamic counterpart PVG [4]. From Tab. 4, we can observe that our strategy consistently reduces the FID by a large margin, and effectively recovering street lanes and cars (i.e. IoU, AP) that are entirely indiscernible in the baseline methods. Unless stated otherwise, all ablation experiments are conducted on two Waymo sequences.
Unreliability mask
The unreliability mask is designed to guide the generation of video diffusion. From Tab. 5 and Fig. 5(a) we can observe: (i) When all regions are masked, where video diffusion model does not know any information from the in-training Gaussian, leading to inferior results; (ii) When the mask is disabled, the generative model can still refine the rendered results toward the video distribution, but with minor refinement; (iii) Equipped with proposed unreliability mask, our full model achieves superior performance in both quantitative and qualitative ablation.
Iterative refinement
The interval of iterative refinement offers the trade-off between quality and efficiency. When the novel trajectory supervision only generated at iteration $T_{0}$ (see Fig. 5(b)), the quality of the novel view is constrained to the in-training Gaussian model which is not fully optimized. To the other extreme, we are unable to afford the time cost of running the diffusion at every step. From Tab. 6, a small interval 500 leads to 4.8×4.8$\times$4.8 × training time. Hence we finally chose the interval 2,000 to strike the balance.
5 Conclusion
In this paper, we present a novel framework that integrates generative priors into the reconstruction of the driving scene from single-trajectory recorded videos, addressing limitations in extrapolating novel views beyond a recorded trajectory. To reasonably leverage the video generative model to provide consistent novel view supervision, we construct a tailored inverse problem to refine the novel view rendered images and progressively enhance the street scene model. The extensive experiments demonstrate that the proposed method significantly improves the synthesis quality for novel trajectories, both in real and AI-generated scenes. These results pave the new way for driving scene synthesis on free-form trajectories.
Supplementary Material
6 More implementation details
Used sequences
The results in Table 1 of the main text are computed from 12 sequences curated from WOD [26]. The names of all sequences are listed below:
  • segment-10444454289801298640
  • segment-10625026498155904401
  • segment-6242822583398487496
  • segment-8822503619482926605
  • segment-10359308928573410754
  • segment-11450298750351730790
  • segment-12496433400137459534
  • segment-15021599536622641101
  • segment-16767575238225610271
  • segment-17860546506509760757
  • segment-3015436519694987712
  • segment-6637600600814023975
For the reconstruction from generated videos, the videos are generated given the following images from nuScenes [2] validation set:
  • n015-2018-09-25-13-17-43+0800__CAM_FRONT__1537853021662460
  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984433512460
  • n015-2018-09-25-13-17-43+0800__CAM_FRONT__1537852899162460
  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984912912460
  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984834412460
  • n015-2018-10-08-15-44-23+0800__CAM_FRONT__1538984833912460
  • n015-2018-07-18-11-41-49+0800__CAM_FRONT__1531885347362460
  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984560512460
  • n015-2018-10-08-15-36-50+0800__CAM_FRONT__1538984563012460
  • n008-2018-08-30-15-31-50-0400__CAM_FRONT__1535657697012404
  • n008-2018-09-18-14-35-12-0400__CAM_FRONT__1537295844362404
  • n008-2018-08-30-15-31-50-0400__CAM_FRONT__1535657694012404
Metric details
Previous driving scene reconstruction methods [38, 8, 13] typically utilize PSNR, LIPIS and SSIM [31] for quality evaluation. However, these metrics are inapplicable for evaluating free-form trajectories due to the absence of corresponding ground-truth images for novel trajectories. Latest research [40] proposed Novel Trajectory Agent IoU (NTA-IoU) and Novel Trajectory Lane IoU (NTL-IoU) for traffic components evaluation. Accordingly, we compared NTA-IoU and NTL-IoU, as shown in Table 2.
Furthermore, we propose improvements to the evaluation metrics of NTA-IoU and NTL-IoU. For vehicle quality, similar to NTA-IoU, we projected the original 3D vehicle bounding boxes onto the new viewpoints to generate corresponding 2D bounding boxes. IoU lacks the precision granularity needed to understand performance across various intersection over union threshold. Therefore, we calculate the Average Precision (AP[50:95]) which provides a more nuanced assessment of detection accuracy.
For lane quality, NTL-IoU calculates the mIoU between the rendered and the ground truth lanes, averaging both foreground lanes and backgrounds. However, including the background can inflate accuracy scores due to the easy detection of non-lane areas. Therefore, following official lane detection benchmarks [3] we focus specifically on IoU with only the foreground to better evaluate lane recognition performance.
For the evaluation of the capability as driving the scene generator, we mainly assess the realism of synthesized images across different trajectories by measuring their FID with the images at recorded trajectories.
More details
For reconstruction from real data, we adopt regularization loss as introduced in [35], including sky loss, LiDAR depth loss and entropy loss for objects. In our framework, the LiDAR depth loss is used for both recorded trajectory and novel trajectory. In addition, we balance the recorded and generated data with 50% of the training data being generated. For reconstruction from generated videos, we additionally include an entropy loss for the opacity of each Gaussian to prevent potential overfitting.
Warping operation $\phi$ in Eq. 7
$(x_{*},y_{*},d)^{\top}$ $=Kp_{\text{camera}}$ (12)
$(x,y)$ $=(\frac{x_{*}}{d},\frac{y_{*}}{d})$ (13)
7 More ablations
Refine strength
[…]
0.4 0.1122 0.5904 98.31 1.0×1.0$\times$1.0 ×
0.6 0.2092 0.6257 90.30 $\times1$
0.8 0.0996 0.5874 92.26 1.2×1.2$\times$1.2 ×
1.0 0.0773 0.5979 117.91 1.3×1.3$\times$1.3 ×
Table 7: Effect of the noise level. In this table, we evaluate the performance under different refine strengths. The metrics are measured on views at the novel trajectory shifted $3m$ from the recorded trajectory.
When we refine the rendered videos using video diffusion, we also study the effect the refine strength (level of noise added to the images). In Tab. 7, we show the metrics for novel trajectory synthesis (IoU, AP, and FID), along with the total optimization time. With a low refine strength, the refined image adheres to the artifact-heavy rendered image, resulting in minimal improvement. On the other hand, a high refine strength yields better novel views but may not keep fidelity to the original scene. Additionally, higher refine strength requires more denoising steps, increasing the training time. Consequently, a refine strength of 0.6 is found to be the best balance.
Other designs
[…]
w/o pose refinement N/A N/A 116.86
w/o depth prior N/A N/A 98.62
Full model N/A N/A 95.73
w/o progress. 0.1075 0.5959 91.22
w/o warp 0.0652 0.5962 119.24
Full model 0.2092 0.6257 90.30
Table 8: Ablations on other design choices. The metric in the first three lines is measured on the generated video, the others are tested on real scene.
We also ablate other design choices in Tab. 8, including the pose refinement and depth prior in the reconstruction from generative video, as well as progressive increasing shifted length and warped condition. The results show that the final quality degrades without any element. The results indicate that the exclusion of any element leads to a decline in the rendering quality and the stability of optimization.
8 Comparison with video generative model
FID $\downarrow$
±$\pm 0m$$\pm 1m$$\pm 2m$
MagicDrive-t [6] 125.07 130.23 138.31
Vista [7] 97.48 99.19 148.07
Recon-only [35] 41.55 95.32 144.01
Ours 36.54 71.47 119.11
Table 9: Comparison with the video-based driving world model. The FID score is calculated between the recorded videos from the nuScenes dataset [2] and the novel trajectory videos synthesized by each method.
In this section, we compare the novel trajectory video synthesis with representative open-sourced generative-based driving world models [7, 6]. Since these works are primarily trained on nuScenes dataset [2], we still conduct the experiments on 12 sequences from this dataset mentioned in Sec. 6. Different from the experiment in Tab. 3, in this experiment, our method takes ground truth recorded trajectory videos as input. The FID between the recorded videos and the synthesized novel trajectory videos from the different methods are reported in Tab. 9.
Note that it is difficult to precisely control the trajectory for generative model. Though we adjust the action control, it still tend to generate videos in the original trajectory. Although all methods do not have significant artifacts, DriveX still demonstrates a clear advantage in FID metrics. This is because it can effectively avoid hallucinations that may be produced by generative models, making its result better aligned with the ground truth. Another advantage difficult to demonstrate by this metric is that DriveX can achieve precise camera pose control. In contrast, other methods are constrained by the diversity and imbalance of actions existing in public driving datasets, thus making it difficult to achieve accurate free-form trajectory synthesis.
9 More results
Figure 6: More comparison results.
Fig. 6 shows more comparison results. We recommend readers to watch the video in the supplementary material.

### v2

@@ Appendix C Additional ablations
Other designs
We also ablate other design choices in Tab. 8, including progressive increasing shifted length and warped condition. The results show that the final quality degrades without any element, which indicates their helpfulness on the rendering quality and the stability of optimization.
@@
Appendix D Comparison with methods incorporating generative prior
Since our work is directly built upon the reconstruction-based driving simulation methods and primarily aims to address the view extrapolation challenge within them, the comparisons in the main paper are mainly against the reconstruction-based approaches. While we acknowledge some concurrent works [44, 20, 28, 4, 41] are sharing a similar scope with ours, most of them have not released their code, making comprehensive and rigorous comparison with them infeasible. Despite that, we try our best to compare with those methods that allow for a reasonable comparison and report the results in Tab. 2. The results indicate that our method performs better than other alternatives, which is also confirmed by the qualitative results in Fig. 4. Compared to DriveDreamer4D [44], our approach substantially eliminates hallucinations and inconsistencies caused by insufficient constraints. Compared to generation-only approach FreeVS [28], our method maintains continuous 3D scene representation, enabling synthesized views to preserve superior detail fidelity and better consistency across frames than that only relying on sparse LiDAR signal. Moreover, our method has significantly faster synthesis speed (34 FPS vs. 0.43 FPS).
Appendix E Application on generated videos
Another intriguing application of the proposed method is transforming AI-generated video into a re-enactable driving world. However, the AI-generated videos present more severe challenges because of the lack of corresponding LiDAR depth and accurate pixel-to-pixel matching between different frames. These challenges further necessitate the novel view supervision provided by the generative priors. Therefore, we devise an additional recipe for the robust optimization of the AI-generated videos.
Details
Specifically, since the generated videos cannot be precisely aligned with the conditioned camera parameters, we employ a feed-forward approach [42] to estimate the camera parameters and corresponding depths for each frame. For scenes containing moving vehicles, we manually annotate their tracklets as the initialization for the local coordinate frames. During optimization, we fine-tune the camera pose to mitigate potential prediction errors. In the computation of unreliability, as the generated video is monocular, we use the nearest previous frames that can include the current shifted view content as the auxiliary source images in Eq. (9) of main paper to ensure well-defined unreliability scores for all areas of the novel view. Besides, we include an entropy loss for the opacity of each Gaussian to prevent potential overfitting.
Results
@@
To quantitatively evaluate this capability, we compared DriveX and the reconstruction-only baseline [37] on 12 driving videos generated by Vista [6]. Due to the evident stylistic gap between the generated videos and those in WOD, we only use the diffusion model trained on general-purpose 3D datasets as generative prior without LiDAR projection as condition. The FID of videos rendered across different novel trajectories is reported in Tab. 9. The results indicate that the realism of rendered videos is effectively preserved within a certain range of shifted length. Compared to the reconstruction-only baseline, our DriveX achieves 4.1% and 2.6% FID reduction in shifting length of ±$\pm 1m$ and ±$\pm 2m$ respectively. To obtain real-world distance, we align the predicted depth and LiDAR depth of the first frame by the least square estimation. These advancements are also clearly illustrated in Fig. 7. As can be seen, the baseline method exhibits significant degradation in views that are far from the recorded trajectory.
Figure 8: Results on the PandaSet dataset [35]. The results are on the novel trajectory shifted $2.0m$ from the recoded trajectory.
Appendix F Results on the PandaSet dataset
To further validate the versatility of the proposed method on a broader range of scene types and sensor configurations, we apply our approach to two sequences indexed $001$ and $003$ from the PandaSet dataset [35], each consisting of 80 frames. The image resolution is downsampled to $450\times 800$ for both training and evaluation. In Fig. 8, we compared our method with the state-of-the-art reconstruction baseline StreetGaussian [37] on the trajectory with shift length $2.0m$. The result suggests that our method can significantly reduce artifacts in synthesized images and improve the discernibility of safety-critical traffic elements, which is also corroborated by our better FID.
Appendix G More quanlitative results
Figure 9: Results on more aggressive view change.
Figure 10: Failure cases. cannot reconstruct a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views.
More aggressive view change.
We test the proposed method on some challenging cases to assess its robustness. In Fig. 9 (left), the view is rendered from the camera elevated by $3m$ above the recorded trajectory, with a pitch angle of $30^{\circ}$. In Fig. 9 (right), we increase the lateral shift length to $6m$. To accommodate these aggressively changed views, we include them in the refined trajectory during optimization. As shown in the results, the baseline method almost totally fails to synthesize meaningful images under these viewpoints, while our solution substantially improves the fidelity.
Failure case.
Despite the robustness and generalization ability on diverse scenes as demonstrated above, it should be noted that the proposed framework only aims to provide regularization for unconstrained reconstruction. As a reconstruction pipeline, it does not work well on totally unseen regions of the scene. For example, in Fig. 10, our method cannot create a feasible appearance from scratch for the side of a vehicle that is invisible from all recorded views and thus lacks proper geometry initialization before the refinement stage.
Figure 11: The evolution of unreliability mask during optimization.
The evolution of unreliability mask during optimization.
In Fig. 11, we present novel view renderings and the corresponding unreliability masks at different training iterations. The results clearly show that the reliable regions gradually expand during iterative refinement, demonstrating the effectiveness of the proposed unreliability mask mechanism–it can serve as a reliable cue for the rendered images rather than introducing compounding errors. Additionally, the distribution of the mask indicates that unreliable regions mainly appear near edge areas of the visual elements, which is the rationale behind the adoption of edge-aware inpainting.
More comparisons with sota methods The video comparison with the state-of-the-art alternates can be found in our project page.